OpenStudy (anonymous):

If 18 meters of high-resistance wire with a diameter of 1.2 mm has a resistance of 10 ohms, what is the resistance of 27 m of the same type of wire with a diameter of 1.5 mm?

OpenStudy (kropot72):

The resistance is proportional to the length and inversely proportional to the diameter of the wire. Therefore the required resistance is given by: \[10\times\frac{27}{18}\times(\frac{1.2}{1.5})^{2}=(you\ can\ calculate)\ \Omega\]